# SynCL: A Synergistic Training Strategy with Instance-Aware Contrastive Learning for End-to-End Multi-Camera 3D Tracking

### v2

@@ 5 Conclusion
In this paper, we delve into the optimization difficulties inherent in joint detection and tracking training in end-to-end trackers. We reveal that the self-attention mechanism hinders multi-task learning through over-deduplicating object queries and introducing self-centric attention on track queries. Based on these analyses, we propose to synergistically optimize multi-task learning through hybrid matching for both queries with dynamic filtering and instance-aware contrastive learning within a weight-shared parallel decoder without self-attention. On the nuScenes dataset, SynCL consistently and significantly outperforms strong baselines, achieving new state-of-the-art performance in multi-camera 3D MOT.

### v3

@@ 5 Conclusion
In this paper, we have delved into the optimization difficulties of joint detection and tracking training in end-to-end 3D trackers. We reveal that the self-attention mechanism in the transformer-based decoder hinders multi-task learning through over-deduplicating object queries and introducing self-centric attention to track queries. Based on these analyses, we have proposed to synergistically optimize multi-task learning through hybrid matching for both queries with dynamic filtering and contrastive learning within a weight-shared parallel decoder without self-attention. Both quantitative and qualitative results on the nuScenes dataset demonstrate the superiority of SynCL with consistent gains, achieving new state-of-the-art performance for the multi-camera 3D MOT task.